A random sample of 16 observations is to be drawn from a normal distribution having mean 11 and standard deviation 3. Let  denote the sample mean. Find, correct to three decimal places

(i)                The probability that  will have a value between 9.2 and 12.2

(ii)              The value of c for which


1
Expert's answer
2021-11-23T17:58:16-0500

Let "x" denote the sample mean: "x\sim N(\mu, \sigma/\sqrt{n})."

Given "\mu=11, \sigma=3, n=16"

(i)


"P(9.2<x<12.2)=P(x<12.2)-P(x\leq 9.2)"

"=P(z<\dfrac{12.2-11}{3/\sqrt{16}})-P(z\leq\dfrac{9.2-11}{3/\sqrt{16}})"

"=P(z<1.6)-P(z\leq -2.4)"

"\approx0.0.9379452007-0.0081975=0.937003"

(ii)


"P(x<c)=0.03"

"P(z<\dfrac{x-11}{3/\sqrt{16}})=0.03"

"c=\dfrac{x-11}{0.75}\approx-1.880794"

"x\approx11-0.75(1.880794)"

"x\approx9.59"